# YOLO Evolution: A Comprehensive Benchmark and Architectural Review of YOLOv12, YOLO11, and Their Previous Versions

### v3

@@ IV-B Discussion
YOLOv3u tiny 33 1 7 33
YOLOv3u 25 28 33 13
YOLOv5un 32 6 3 5
YOLOv5us 29 8 10 11
YOLOv5um 21 15 15 21
YOLOv5ul 18 21 25 27
YOLOv5ux 22 31 31 32
YOLOv8n 31 5 5 6
YOLOv8s 26 9 13 12
YOLOv8m 19 18 19 24
YOLOv8l 17 23 26 26
YOLOv8x 11 30 32 31
YOLOv9t 23 14 4 1
YOLOv9s 10 16 12 7
YOLOv9m 5 22 18 18
YOLOv9c 13 26 23 23
YOLOv9e 16 33 28 30
YOLOv10n 30 2 6 4
YOLOv10s 23 3 11 8
YOLOv10m 9 9 14 14
YOLOv10b 11 12 22 17
YOLOv10l 14 16 24 20
YOLOv10x 2 26 27 25
YOLO11n 27 3 1 3
YOLO11s 19 6 8 10
YOLO11m 1 11 17 15
YOLO11l 2 13 20 19
YOLO11x 8 23 29 28
YOLOv12n 27 19 1 2
YOLOv12s 7 25 8 9
YOLOv12m 5 20 16 16
YOLOv12l 4 28 21 22
YOLOv12x 15 32 30 29
The analysis of Table IX yields critical observations:
IV-B1 Accuracy
YOLO11m consistently emerged as a top performer, frequently ranking among the highest, closely followed by YOLOv10x, YOLO11l, YOLOv9m, and YOLO11x. This underscores the robust performance of the YOLO11 family across varying IoU thresholds and object sizes, which can be attributed to their use of C2PSA for the preservation of contextual information, leading to improved convergence and overall performance. The inclusion of two small kernel convolutions and partial self-attention modules in these models helped capture more spatial features, leading to better accuracy in detecting both large and small objects. Additionally, the fine-tuning of anchor boxes and multi-scale training has enhanced the detection of objects at different resolutions, which is crucial in achieving high accuracy across diverse datasets.
Conversely, YOLOv3u-tiny exhibited the lowest accuracy, particularly in the Africa Wildlife and Ships and Vessels datasets. YOLOv5un and YOLOv8n show slightly better but still sub-par results. This suggests that YOLO11 models are currently the most reliable for applications demanding high accuracy.
The YOLOv3u-tiny model’s underperformance can be attributed to its older architecture, which lacks the optimizations found in newer YOLO models. YOLOv3u is based on the Darknet-53 framework, which, although efficient for its time, struggles to capture complex patterns and fine details required for high-accuracy detection. Its relatively shallow architecture and absence of more recent enhancements such as attention mechanisms or advanced feature fusion techniques make it less effective in handling intricate object detection tasks, especially in datasets with a wide variety of object sizes, like those in the Africa Wildlife and Ships and Vessels datasets.
In contrast, the YOLOv9 models show effectiveness in detecting objects across various sizes and IoU thresholds, but they still struggle with small objects, particularly in the Ships and Vessels dataset. The YOLOv9 family benefits from improved training strategies, such as anchor box refinement and better scale handling, but may still fall short in detecting small objects in crowded or cluttered scenes due to its reliance on larger-scale feature maps and a less efficient processing pipeline for fine-grained features.
The YOLOv10 family, despite its later introduction, exhibited relatively lower accuracy in the Traffic Signs and Africa Animals datasets, resulting in an average accuracy drop of 2.075% compared to the YOLOv9 models in those datasets. YOLOv10’s adoption of the One-to-One Head approach instead of Non-Maximum Suppression (NMS) for defining bounding boxes contributed to this decrease. The One-to-One Head approach relies on predicting a single bounding box per object, which can be a limitation in scenarios with overlapping or closely packed objects. NMS, on the other hand, is more effective at filtering out redundant predictions, which can improve detection accuracy in these types of datasets. Therefore, the One-to-One Head’s inability to handle multiple overlapping predictions adequately explains YOLOv10’s underperformance, especially in datasets like Traffic Signs, where multiple signs are often close together.
Although the YOLOv12 family includes strong models like YOLOv12l, which ranked 4th in accuracy, it falls short compared to previous iterations, particularly considering its status as the newest YOLO version. It shows a -2.72% decline compared to YOLOv9 in the African Animals dataset and a -1% decrease in accuracy compared to YOLO11 in the Ships and Vessels dataset. This decline in accuracy is largely due to architectural choices. R-ELAN and area attention, while innovative, compromised global dependency modeling, impacting accuracy. Additionally, the adjusted MLP ratio prioritized certain computational aspects over feature representation depth, limiting its ability to capture intricate patterns. Consequently, YOLOv12 underperformed compared to the more balanced YOLOv11 and YOLOv9 families, demonstrating that newer architectural changes did not yield the expected gains in accuracy.
Similarly, YOLOv8n, though it performed better than YOLOv3u, still showed lower accuracy than the top-performing models in specific datasets. This model demonstrated a good balance between speed and accuracy but struggled with intricate object recognition tasks. The simplicity of its architecture, without more advanced modules like self-attention or multi-scale training, likely held it back from achieving the high accuracy observed in models like YOLO11 and YOLOv9.
Lastly, the YOLO11 models stand out for their ability to consistently achieve the highest accuracy. The use of advanced mechanisms like C2PSA and two kernel convolutions allowed for improved feature extraction across multiple scales, enhancing their ability to detect objects of varying sizes and complexities. These models’ robust performance across a variety of datasets, including challenging ones like Africa Wildlife and Ships and Vessels, demonstrates their ability to handle a wide range of object detection tasks with exceptional accuracy.
IV-B2 Computational Efficiency
YOLOv10n consistently outperformed other models in terms of speed and GFLOPs count, ranking among the top for speed across all datasets and 5th in terms of GFLOPs. Its use of CSPNet as backbone design and optimization strategies allowed for high processing speed without significantly compromising performance. YOLOv3u-tiny, YOLOv10s, and YOLO11n also demonstrated excellent computational efficiency, with their lightweight architectures enabling faster inference times and lower resource usage.
Despite its accuracy advantages, YOLOv12 ranked as the worst family in terms of processing time and second in GFLOPs count. This is largely due to its attention-centric architecture, which enhances accuracy but significantly increases computational complexity. The integration of Area Attention and R-ELAN, while effective in capturing spatial features, introduces substantial overhead in feature extraction and aggregation, leading to slower inference times. Additionally, the reliance on FlashAttention contributes to inefficient memory access patterns, further impacting processing speed. Although these architectural choices improve contextual feature extraction, they hinder computational efficiency, making YOLOv12 less suitable for real-time applications compared to its predecessors. This performance trade-off illustrates that while attention mechanisms can elevate accuracy, they come at the cost of increased latency and computational demands.
YOLOv5un, YOLOv5us, and YOLOv8n exhibited strong efficiency, providing a good balance between speed and accuracy. These models performed well in both inference times and GFLOPs, with YOLOv5un showing slightly better speed than YOLOv8n potentially due to the implementation of C2k (CSP with larger kernel sizes) in YOLOv8 architecture that increases accuracy but with a trade-off in computational efficiency. YOLOv8n, despite its good speed performance, had a larger size and computational cost compared to some other lightweight models.
While YOLOv3u-tiny ranked highly in speed, its performance in accuracy was subpar, particularly in datasets like Africa Wildlife and Ships and Vessels. This highlights a critical trade-off: while speed is crucial for real-time applications, it becomes irrelevant if it severely compromises accuracy. YOLOv3u-tiny’s lightweight architecture, while optimized for faster inference, failed to detect objects accurately across varied and complex datasets, making it unsuitable for applications where precision is key. On the other hand, YOLOv3u struggled in both speed and accuracy, ranking poorly in both categories. Its reliance on the outdated Darknet-53 backbone contributed to this underperformance, as it was unable to meet modern detection challenges. The model’s low ranking in both computational efficiency and accuracy makes it less viable compared to newer architectures, such as YOLOv11 and YOLOv10, which offer better performance across both metrics.
In contrast, YOLO11 models demonstrated significant speed improvements, thanks to the C3k2 block, which reduces computational overhead while maintaining high accuracy. These models were faster than YOLOv10 by 1.41% and YOLOv9 by 31% on average, making them ideal for real-time applications requiring rapid processing without sacrificing accuracy. While YOLOv9 models excelled in accuracy, their slower inference times made them less ideal for time-sensitive tasks. YOLOv10, slightly slower than YOLO11, offered a good balance between speed and accuracy, making it well-suited for real-time applications.
Lastly, YOLOv5 and YOLOv8 families also performed well, with YOLOv5un ranking near the top for speed and efficiency, making it a solid choice for applications where both quick processing and solid accuracy are required.
IV-B3 Model Size
YOLOv9t emerged as the smallest model, ranking first across all three datasets in terms of model size thanks to its use of GELAN architecture leading to more efficient use of parameters. It is followed closely by YOLOv12n, YOLO11n, and YOLOv10n. Specifically, YOLOv10n benefits from the Spatial-Channel Decoupled Downsampling technique, which effectively reduces the model size while maintaining high accuracy, making it a highly efficient choice for resource-constrained environments. This technique allows for better utilization of parameters, enabling smaller models to perform similarly to larger counterparts.
In contrast, YOLOv3u was the largest model, ranking poorly in terms of both size and accuracy. This inefficiency stems from its reliance on the outdated Darknet-53 architecture, which is larger and more computationally expensive compared to more modern designs. The size of YOLOv3u further exacerbates its subpar performance, as the model requires more resources and longer inference times, without corresponding improvements in detection accuracy. Its larger size makes it less suitable for deployment in environments where computational resources are limited or where real-time processing is essential.
This comparison highlights the trend in newer YOLO versions, such as YOLOv9, YOLO10, and YOLO11, where optimizations have led to smaller, more efficient models without significant sacrifices in accuracy or speed. These advancements make them more attractive for modern applications, where both model size and computational efficiency are critical. However, despite being the latest addition, YOLOv12 did not follow this trend, ranking third in terms of model size and computational efficiency.
IV-B4 Overall Performance
Considering accuracy, speed, size, and GFLOPs, YOLO11m, YOLO11n, YOLO11s, and YOLOv10s emerged as the most consistent performers. These models demonstrated a strong balance between high accuracy, low processing time, and efficient disk usage, making them suitable for a wide range of applications where both speed and accuracy are crucial. YOLO11 models, in particular, benefited from recent enhancements such as the C3k2 block and C2PSA module, which contributed to their superior performance in these areas. The incorporation of these innovations allowed the YOLO11 models to handle larger datasets more effectively without compromising on speed or efficiency.
On the other hand, YOLOv9e, YOLOv5ux, and YOLOv3u showed poor results across all metrics. These models were computationally inefficient, struggling with high GFLOPs counts and slower processing times. YOLOv3u, in particular, ranked poorly due to its reliance on the outdated Darknet-53 architecture, which is no longer optimized for modern object detection tasks. Despite its larger size, it failed to offer the necessary accuracy, making it less suitable for practical applications. Similarly, YOLOv5ux and YOLOv9e showed poor performance in terms of both speed and accuracy, which further limits their real-world usability.
YOLO11 models were clearly the best performers overall, owing to their modern design and architectural improvements, making them highly suitable for a wide range of use cases. YOLOv10 models, although slightly underperforming in accuracy, excelled in computational efficiency due to their use of the One-to-One head for predictions. This design allowed YOLOv10 to process data more quickly and efficiently, making it ideal for time-sensitive applications where speed is prioritized over perfect accuracy.
While YOLOv9 models showed underperformance in terms of computational efficiency, they still remain competitive with YOLOv10 and YOLO11 in terms of accuracy, particularly due to their PGI integration, which helps in enhancing detection precision. This makes YOLOv9 a viable option for applications where high accuracy is prioritized over speed.
Concerning the YOLOv12 family, it showcased mediocre results, underperforming compared to its predecessors, particularly YOLOv9 and YOLO11, in both accuracy and speed. Despite introducing innovative components like the Area Attention Module (A2), Residual Efficient Layer Aggregation Networks (R-ELAN), and FlashAttention, these architectural changes contributed to increased computational overhead and slower inference times. The Area Attention Module segmented the feature map to maintain a large receptive field but limited contextual understanding, leading to lower accuracy. Meanwhile, R-ELAN improved feature aggregation but introduced latency due to its complex residual connections. FlashAttention aimed to optimize memory access but added processing overhead, further slowing down the model. Although YOLOv12 leveraged convolution operators for efficiency, it failed to offset the complexity introduced by its attention mechanisms. Overall, YOLOv12 struggled to balance speed and accuracy, making it less favorable compared to the more efficient and accurate YOLOv9, YOLOv10, and YOLO11 models.
In addition, YOLOv8 and YOLOv5u achieved competitive results, surpassing YOLOv3u in terms of accuracy. This improvement can be attributed to their more recent architectures, which allowed them to handle modern detection tasks better. However, despite their higher accuracy compared to YOLOv3u, they still fell short when compared to newer models like YOLOv9, YOLOv10, and YOLO11 in both accuracy and overall performance. While YOLOv8 and YOLOv5u had faster processing times than YOLOv9, their overall performance was still inferior to that of the YOLO11 and YOLOv10 models, which maintained a better balance between speed and accuracy.
IV-B5 The Rise of YOLO11 Over YOLOv8
Although YOLOv8 [27] has been the algorithm of choice for its versatility in tasks such as pose estimation, instance segmentation, and oriented object detection (OBB), YOLO11 [28] has now emerged as a more efficient and accurate alternative. With its ability to handle the same tasks while offering improved contextual understanding and better architectural modules, YOLO11 sets a new standard in performance, surpassing YOLOv8 in both speed and accuracy across various applications.
This paper contributes by providing empirical evidence of YOLO11’s superior performance in both accuracy and computational efficiency. Our findings establish YOLO11 as a clear advancement in the YOLO family, setting a new benchmark for object detection. By incorporating novel modules, YOLO11 addresses both the need for high accuracy and low-latency processing, offering valuable insights for deploying YOLO-based models in practical, real-world applications.
IV-B6 YOLOv9: Efficient Yet Slow
Although YOLOv9 [70] excels in GFLOPs count and model size, outperforming many other models, it paradoxically ranks as one of the slowest in terms of inference speed. This unusual combination of high computational efficiency but low speed can be attributed to YOLOv9’s architectural choices, such as the use of Programmable Gradient Information (PGI) and Generalized Efficient Layer Aggregation Network (GELAN), which optimize computational performance but add significant latency during inference.
This paper contributes by highlighting the trade-offs in YOLOv9’s design, showing that while its smaller model size and lower GFLOPs count make it ideal for resource-constrained environments, its slow inference times make it unsuitable for time-sensitive applications. Our findings suggest that YOLOv9’s performance is best suited for scenarios where computational efficiency is prioritized over speed, offering valuable insights for selecting models based on specific application needs.
IV-B7 YOLOv12: Ambitious but Underwhelming
YOLOv12, despite being the latest addition to the YOLO family, delivered subpar results, failing to outperform its predecessors, particularly YOLOv9, YOLOv10, and YOLO11, in both accuracy and speed. This underperformance is primarily due to its complex architectural choices, including the Area Attention Module (A2), Residual Efficient Layer Aggregation Networks (R-ELAN), and FlashAttention, which introduced significant computational overhead. Although these modules aimed to enhance contextual understanding and memory efficiency, they ultimately compromised speed and accuracy due to increased latency and reduced spatial awareness. This work contributes by revealing the trade-offs inherent in YOLOv12’s ambitious design, demonstrating that its added complexity did not translate into superior performance.
IV-B8 Dataset Size
@@
IV-B9 Impact of Training Datasets
@@
The performance of YOLO models is influenced by the training datasets used, as shown in Tables VI, VII, and VIII. Different datasets yield varying results and top performers, indicating that dataset complexity affects algorithm performance. This underscores the importance of using diverse datasets during benchmarking to obtain comprehensive results on the strengths and limitations of each model.
This discussion highlights the need for a balanced consideration of accuracy, speed, and model size when selecting YOLO models for specific applications. The consistent performance of YOLO11 models across various metrics makes them highly recommended for versatile situations where accuracy and speed are essential. Meanwhile, YOLOv10 models can perform similarly while achieving faster processing times and with smaller model sizes. Additionally, YOLOv9 can deliver comparable results in terms of accuracy but sacrifices speed, making it suitable for applications where precision is prioritized over rapid processing.
V Conclusion

### v4

@@ 5 Discussion
YOLOv3u 7 1 4 7
YOLOv5u 6 4 4 5
YOLOv8 5 5 7 6
YOLOv9 3 6 6 4
YOLOv10 4 2 3 1
YOLO11 1 3 1 2
YOLOv12 1 7 2 3
5.1 Performance Analysis by Model Scale
Nano Models
Among nano models, YOLOv9t ranked highest in accuracy and speed, benefiting from its lightweight GELAN architecture, which optimizes parameter efficiency while maintaining strong detection capabilities. However, YOLOv10n, despite its computational efficiency, exhibited lower accuracy due to its reliance on the One-to-One head and NMS-Free training, which struggled with densely packed objects.
Small Models
YOLOv12s emerged as the best small model in terms of accuracy, likely due to its R-ELAN-based feature aggregation. However, it performed poorly in speed, as the added complexity of Area Attention (A2) increased inference time. In contrast, YOLOv10s and YOLO11s provided a balanced performance, with YOLO11s leveraging the C3k2 block for improved feature extraction and YOLOv10s excelling in speed due to its efficient head design.
Medium Models
YOLO11m achieved the best balance of accuracy and efficiency in this category, owing to C2PSA, which enhances spatial feature capture. YOLOv10m outperformed all models in terms of computational efficiency but had lower accuracy due to its reliance on NMS-Free training. Meanwhile, YOLOv9m, while achieving strong accuracy, ranked lower in efficiency due to its reliance on gradient-based optimizations, which added computational overhead.
Large Models
YOLO11l outperformed other large models in accuracy and speed, thanks to the integration of C2PSA and refined convolutional operations. In contrast, YOLOv9c and YOLOv12l struggled with efficiency, with YOLOv9c being computationally expensive and YOLOv12l suffering from latency due to FlashAttention’s memory overhead.
Extra-Large Models
YOLO11x provided the best balance of accuracy and computational efficiency, with its C3k2-based optimization reducing inference overhead. However, YOLOv9e and YOLOv12x exhibited slower speeds and inefficient GFLOPs utilization, highlighting the limitations of their respective architectures when scaling up.
5.2 Performance Analysis by Family
YOLOv12
YOLOv12 delivered strong accuracy results, ranking first alongside YOLO11 in accuracy across models. This performance can be attributed to its integration of the Area Attention Module (A2) and Residual Efficient Layer Aggregation Networks (R-ELAN), which enhanced feature extraction and contextual understanding. However, this came at the cost of computational efficiency, as the added complexity introduced significant latency and reduced processing speed. FlashAttention, while designed to optimize memory access, further contributed to increased inference time. While YOLOv12 excelled in accuracy, its slow processing speed and higher computational demands limited its overall practicality, highlighting the trade-off between architectural complexity and real-world efficiency.
YOLO11
The YOLO11 family consistently ranked among the best due to its C3k2 block and C2PSA, enhancing efficiency and contextual understanding by ranking first in terms of accuracy and GFLOPs count and among the first in terms of speed and size. These innovations enabled YOLO11 to achieve high accuracy with low computational overhead, making it a well-balanced choice. While YOLOv8 was widely used for various tasks such as Pose Estimation and OBB, YOLO11 has emerged as a superior alternative, offering better feature extraction, inference speed, and accuracy. This study establishes YOLO11 as a new benchmark, demonstrating its high detection accuracy and low-latency processing, and various capabilities, making it ideal for real-world applications requiring both speed and precision.
YOLOv10
YOLOv10 models were highly efficient, particularly in speed and size, due to the One-to-One head and NMS-Free training. However, their accuracy was subpar compared to YOLOv9, YOLO11, and YOLOv12, particularly in scenarios requiring precise bounding box refinement, such as the Traffic Signs dataset.
YOLOv9
Despite strong accuracy due to PGI and GELAN, YOLOv9 models suffered from slow inference times, limiting their real-time applicability. Their architectural choices prioritized precision over efficiency, making them suitable for applications where accuracy is paramount but speed is less critical.
YOLOv8, YOLOv5u, and YOLOv3u
While YOLOv8, YOLOv5usurpassed YOLOv3u in accuracy, they lagged behind newer models in overall efficiency. YOLOv8 introduced the C2k module for improved feature extraction, but its higher computational cost made it less efficient than more recent architectures. With YOLO11 offering superior accuracy, speed, and efficiency, YOLOv8 is no longer the go-to model for real-world applications, as YOLO11 provides a better balance between performance and computational overhead. Meanwhile, YOLOv5u improved on YOLOv3u but lacked modern enhancements like self-attention, limiting its scalability.
Overall, the rankings highlight the dominance of YOLO11 as the most balanced family, with YOLOv10 excelling in efficiency and YOLOv9 maintaining strong accuracy at the cost of speed. YOLOv12, despite its ambitious design and exceptional accuracy, failed to outperform its predecessors due to increased latency, reinforcing the importance of balancing architectural complexity with real-world feasibility.
5.3 Dataset Size
@@
5.4 Impact of Training Datasets
@@
The performance of YOLO models is influenced by the training datasets used, as shown in Figures 12, 16, and 20. Different datasets yield varying results and top performers, indicating that dataset complexity affects algorithm performance. This underscores the importance of using diverse datasets during benchmarking to obtain comprehensive results on the strengths and limitations of each model. Additionally, this highlights the need for a balanced consideration of accuracy, speed, and model size when selecting YOLO models for specific applications.
6 Real-Life Applications
Table 5: YOLO Model Recommendations for Real-Life Applications
Scenario Recommended YOLO Models Application Examples
Computationally Constrained Environments YOLOv10(n/s/m) Drone surveillance, embedded systems, battery-powered devices [29, 80, 25, 27, 9, 8]
YOLO11(n/s/m)
Real-Time Monitoring and Rapid Response YOLOv10(n/s/m) Autonomous vehicles, disaster response, traffic monitoring [29, 80, 25, 75, 27, 9, 8]
YOLO11(n/s/m)
Detection of Small Objects YOLO11(m/l/x) Wildlife tracking, satellite imaging of ships and small vehicles [75, 27, 9, 8]
YOLOv12(m/l/x)
Detection of Large Objects YOLOv8(s/m/l/x) Satellite imaging (buildings, forests, urban development) [83, 55, 43]
YOLOv9(t/s/m/c/e)
Overlapping and Densely
Packed Objects
YOLOv9(t/s/m/c/e)
Traffic intersections,
wildlife in dense foliage
Oriented and Rotated Object Detection YOLO11 (OBB Support) Satellite imagery, aerial monitoring of ships [46, 21]
YOLOv12 (OBB Support)
Handling Large Datasets YOLO11(m/l/x) City-wide monitoring, biodiversity mapping [5, 76]
YOLOv12(m/l/x)
Handling Small Datasets YOLOv9(t/s/m/c)
Anti-poaching monitoring,
niche industrial tasks [1, 64, 84]
The advancements in YOLO models have enabled their application across diverse real-world scenarios, each with distinct demands for model performance, such as computational efficiency, accuracy, and adaptability to varying object complexities. Based on the benchmark of various YOLO models and related papers, this section highlights how different models excel under specific scenarios, making them suitable for deployment in various tasks across industries, as shown in Table 5.
6.1 Computationally Constrained Environments
In scenarios with limited computational resources, such as drones, embedded systems, and battery-powered devices, models must be lightweight and efficient to avoid draining power or overloading hardware. The nano, small, and medium-sized models of the YOLO11 and YOLOv10 families are optimal choices due to their high accuracy, compact model sizes, low inference times, and low GFLOPs count leading to minimal resource consumption, as also seen in related papers [29, 80, 25, 27, 9, 8].
6.2 Real-Time Monitoring and Rapid Response
For applications requiring real-time detection, such as autonomous driving, surveillance, and disaster response, models need to balance accuracy and speed. As illustrated by this study and related research papers [29, 80, 25, 75, 27, 9, 8], the nano, small, and medium-sized YOLO11 and YOLOv10 models demonstrate strong performance in maintaining fast inference while accurately detecting objects across various scales.
6.3 Detection of Small and Large Objects
Object detection models often face the challenge of accurately identifying small and large objects. Small objects, such as wildlife or distant vehicles, require models capable of capturing fine spatial details. Medium, large, and extra-large-sized YOLO11 and YOLOv12 models excel in these scenarios, as illustrated in our paper and related work [75, 27, 9, 8]. Meanwhile, all YOLOv9 and YOLOv8 models offer robust performance for detecting large objects, making them ideal for scenarios such as satellite imaging of buildings, forests, and urban development [83, 55, 43].
6.4 Overlapping and Densely Packed Objects
Detecting objects that overlap or appear densely packed is a major challenge in traffic intersections or wildlife monitoring scenarios. In these cases, models must efficiently distinguish between objects and avoid false positives. All YOLOv9 models are highly effective in handling overlapping objects, as seen in this benchmark and related work [46, 21]. This suits them, particularly for applications like detecting animals in dense foliage or distinguishing closely packed traffic signs.
6.5 Oriented Object Detection
Objects often appear at varying angles or orientations in satellite imaging and certain surveillance applications, requiring models with specialized detection capabilities. The YOLO11 and YOLOv12 families’ support for OBB enables effective detection of rotated objects, achieving higher accuracy than YOLOv8 models.
6.6 Handling Large and Very Small Datasets
Large-scale datasets with diverse object categories require models that generalize well without overfitting. Medium, large, and extra-large YOLO11 and YOLOv12 models excel in such scenarios, as showcased in our paper and related studies [5, 76]. For tiny datasets, where overfitting is a concern, all YOLOv9 models are optimal choices. Their compact sizes and efficient parameter usage enable effective learning from limited data [1, 64, 84].
7 Conclusion